OpenStudy (anonymous):

solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair. (x'y) 3x-2y=9 -3x+2y=-9

OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 3x - 2y = 9 and -3x + 2y = -9 the 2nd equation is just the 1st multiplied by -1 it will be infinitely many solutions

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

np

OpenStudy (anonymous):

y = (3 (x-3))/2